# LibIQ: Toward Real-Time Spectrum Classification in O-RAN dApps

### v2

@@ Abstract
Index Terms:
Open RAN, dApps, Real-Time Control Loops, Spectrum Classification
@@ I Introduction
The Open Radio Access Network (RAN) paradigm is gaining momentum in the cellular networks thanks to its open and modular approach that promotes interoperability and innovation by disaggregating the RAN functionalities over white-box hardware and different software components [1]. In the O-RAN architecture, network control is realized through closed control loops—continuous cycles of data collection and control actions, in which specialized applications can adapt the network to evolving conditions. However, in the current standard defined by the O-RAN ALLIANCE, two key limitations persist: the lack of attention for the user-plane and physical-layer data such as I/Q samples, and the support for sub-10 ms control loops, ruling out the real-time feedback needed for advanced low-latency use cases, such as beamforming, spectrum sharing, and more [2]. The introduction of real-time control loops will enable direct action at the lower layers of the RAN, allowing adaptive responses to radio conditions while minimizing delays. Moreover, a direct access to I/Q samples enables classification systems to detect anomalous signals and interference, enhancing attack detection, dynamic resource management, and spectrum sharing to protect network integrity of the 5th generation (5G) and beyond cellular networks. To address these challenges, dApps [3] are introduced as software components deployed alongside Central Unit (CU) and Distributed Unit (DU), where user-plane and I/Q data is directly accessible, enabling Artificial Intelligence (AI)/Machine Learning (ML) routines to operate in real-time within the RAN. However, the challenges of processing large I/Q data in real-time remains, requiring specialized AI/ML models optimized for edge inference [3].
@@
To validate the effectiveness of our approach, we conduct experiments in two different Software-defined Radio (SDR)-based environments: the Colosseum network emulator [4], and an Over-The-Air (OTA) testbed. The main contributions of this work are:
@@ II-A O-RAN and dApps
The O-RAN paradigm is a new architectural concept for cellular networks based on disaggregation and programmability of the RAN functions for enabling a multi-vendor, data-driven approach [1]. O-RAN disaggregates the classic Next Generation Node Base (gNB) into CU, DU, and Radio Unit (RU), introducing the RAN Intelligent Controllers (RICs) as the entity responsible for the control and managements of the units through open interfaces. Thanks to this approach, the RICs become centralized abstractions of the networks that can manage near-real-time and non-real-time tasks through xApps and rApps. However, at the current time, there is no standardized way to access real-time, i.e., sub-10 ms timescales, parameters in the RAN, therefore no real-time control is available. To address these issues, a recent research report [2] from the O-RAN next Generation Research Group (nGRG) explores the concept of dApps. dApps are microservices co-located with CU and DU that enable the real-time exposure and processing of RAN data that otherwise would not be available to the RICs due to latency or privacy constraints [3]. In this paper, we leverage dApps by embedding our library’s CNN to classify RF signal types based on processed user-plane data. To the best of our knowledge, this is the first work to apply the dApp framework to perform real-time spectrum classification within the O-RAN context.
@@ II-B Spectrum Classification with Deep Learning
For instance, [5] applies a residual CNN to normalized power spectrum data for binary spectrum sensing, effectively distinguishing between signal-plus-noise and noise-only scenarios. Meanwhile, [6] integrates a CNN with an Long Short Term Memory (LSTM) to analyze STFT-based spectrograms, achieving high accuracy in identifying wireless technologies such as Wi-Fi and 5G. Other works focus on CNN-based architectures specifically optimized for hardware efficiency. The authors in [7] employ a three-layer CNN designed to minimize computational complexity, making it suitable for embedded systems with constrained processing power. In contrast, [8] transforms I/Q samples into image representations and applies the YOLOv8x object detection model, achieving 77.78% accuracy on a small dataset. Similarly, [9] finds that frequency-domain I/Q samples provide better inputs for CNN-based modulation classification compared to time-domain representations, underscoring the importance of input feature selection.
To improve latency and efficiency, alternative CNN architectures avoid spectrogram conversion. Spectrum Stitching [10] uses a U-Net with a non-local block for direct I/Q processing, reducing computational overhead and speeding up inference. While this method eliminates feature extraction, it requires extensive training to generalize across signal variations, balancing accuracy and efficiency. In the context of O-RAN-based implementations, ChARM [11] provides a real-time, data-driven framework for dynamic spectrum sharing in O-RAN networks, enabling spectrum sensing, interference detection. However, its integration with O-RAN remains unclear, and its classifier handles a limited set of RF signals.
Our approach differs from previous works in multiple aspects. Unlike spectrogram-based classification ([6][8]) or hardware-efficient CNNs ([7]), we process frequency-domain I/Q samples directly using a CNN trained on time-series data that performs multinomial classification, in contrast with the binary one of [5]. Before classification, we apply an Energy Peak Detector to isolate relevant signals and reduce dimensionality, optimizing both performance and efficiency. In terms of results, our method achieves 97.8% classification accuracy, significantly outperforming image-based classification ([8], 77.78%) and offering superior real-time adaptability compared to methods based on static spectrograms ([9], approx 90%). Furthermore, our spectrum classification model is the only one that has been tested within a real dApp framework, interacting directly with an operating gNB, unlike previous works that primarily rely on offline datasets or simulated environments or substantial modification of the O-RAN architecture [11].
@@ IV-A Dataset collection and preprocessing
We leverage and extend the dApp framework of [3] to access and collect the I/Q samples in real-time. The dApp is designed with a sensing periodicity of approximately 8 ms, meaning it periodically provides data to LibIQ for labeling. Once collected, the I/Q samples are organized into complex input vectors, each containing real and imaginary components. Each vector consists of 1,536 I/Q samples, representing data collected from spectrum sensing. These vectors are stored in a binary file, with each file containing 100 input vectors.
We collected data from two different environments: one in an emulated environment using Colosseum [4], and the other in a real environment using transmissions over the air. We use a 5G standalone (SA) deployment of one gNB with OpenAirInterface (OAI), operating at a center frequency of 3.6192 GHz with a 40 MHz bandwidth that onboard our dApp. As shown in Figure 1, for the OTA testbed we connect a Commercial Off-the-Shelf (COTS) One Plus Nord phone to the network to generate downlink traffic even though in this work we focus more on the external than the User Equipment (UE) performance. To generate external signals, we employed GNU Radio, the siggen utility of the uhd library for SDRs, and the SCOPE framework [12], each transmitting different signal types, as shown in Table I. The center frequency of the transmitted signals varied among 3.6042 GHz, 3.6142 GHz, 3.6242 GHz, and 3.6342 GHz, while the sampling rate remained fixed at 1 MHz throughout all transmissions.
@@
Sine $90$ $25$ $1$ $1$ Radar Siggen
Triangular $100$ $25$ $1$ $1$ Triangular GNU Radio
Square $100$ $25$ $1$ $1$ Square GNU Radio
Uniform $80$ $20$ $1$ $1$ Jammer Siggen
LTE /// $20$ /// $0.9$ LTE SCOPE [12]
@@ V Results
Latency [ms] Time required for a single prediction $2.23$$\pm 0.48$ $1.13$$\pm 0.04$ $0.92$$\pm 0.11$ $0.90$$\pm 0.05$

### v3

@@ I Introduction
The Open Radio Access Network (RAN) paradigm is gaining momentum in the cellular networks thanks to its open and modular approach that promotes interoperability and innovation by disaggregating the RAN functionalities over white-box hardware and different software components [polese2023understanding]. In the O-RAN architecture, network control is realized through closed control loops—continuous cycles of data collection and control actions, in which specialized applications can adapt the network to evolving conditions. However, in the current standard defined by the O-RAN ALLIANCE, two key limitations persist: the lack of attention for the user-plane and physical-layer data such as I/Q samples, and the support for sub-10 ms control loops, ruling out the real-time feedback needed for advanced low-latency use cases, such as beamforming, spectrum sharing, and more [dappsOranReport]. The introduction of real-time control loops will enable direct action at the lower layers of the RAN, allowing adaptive responses to radio conditions while minimizing delays. Moreover, a direct access to I/Q samples enables classification systems to detect anomalous signals and interference, enhancing attack detection, dynamic resource management, and spectrum sharing to protect network integrity of the 5th generation (5G) and beyond cellular networks. To address these challenges, dApps [lacava2025dapps] are introduced as software components deployed alongside Central Unit (CU) and Distributed Unit (DU), where user-plane and I/Q data is directly accessible, enabling Artificial Intelligence (AI)/Machine Learning (ML) routines to operate in real-time within the RAN. However, the challenges of processing large I/Q data in real-time remains, requiring specialized AI/ML models optimized for edge inference [lacava2025dapps].
@@
To validate the effectiveness of our approach, we conduct experiments in two different Software-defined Radio (SDR)-based environments: the Colosseum network emulator [polese2024colosseum], and an Over-The-Air (OTA) testbed. The main contributions of this work are:
@@ II-A O-RAN and dApps
The O-RAN paradigm is a new architectural concept for cellular networks based on disaggregation and programmability of the RAN functions for enabling a multi-vendor, data-driven approach [polese2023understanding]. O-RAN disaggregates the classic Next Generation Node Base (gNB) into CU, DU, and Radio Unit (RU), introducing the RAN Intelligent Controllers (RICs) as the entity responsible for the control and managements of the units through open interfaces. Thanks to this approach, the RICs become centralized abstractions of the networks that can manage near-real-time and non-real-time tasks through xApps and rApps. However, at the current time, there is no standardized way to access real-time, i.e., sub-10 ms timescales, parameters in the RAN, therefore no real-time control is available. To address these issues, a recent research report [dappsOranReport] from the O-RAN next Generation Research Group (nGRG) explores the concept of dApps. dApps are microservices co-located with CU and DU that enable the real-time exposure and processing of RAN data that otherwise would not be available to the RICs due to latency or privacy constraints [lacava2025dapps]. In this paper, we leverage dApps by embedding our library’s CNN to classify RF signal types based on processed user-plane data. To the best of our knowledge, this is the first work to apply the dApp framework to perform real-time spectrum classification within the O-RAN context.
@@ II-B Spectrum Classification with Deep Learning
For instance, [zheng2020spectrum] applies a residual CNN to normalized power spectrum data for binary spectrum sensing, effectively distinguishing between signal-plus-noise and noise-only scenarios. Meanwhile, [zhang2021signal] integrates a CNN with an Long Short Term Memory (LSTM) to analyze STFT-based spectrograms, achieving high accuracy in identifying wireless technologies such as Wi-Fi and 5G. Other works focus on CNN-based architectures specifically optimized for hardware efficiency. The authors in [lees2019deep] employ a three-layer CNN designed to minimize computational complexity, making it suitable for embedded systems with constrained processing power. In contrast, [amiciradiodl] transforms I/Q samples into image representations and applies the YOLOv8x object detection model, achieving 77.78% accuracy on a small dataset. Similarly, [hauser2017signal] finds that frequency-domain I/Q samples provide better inputs for CNN-based modulation classification compared to time-domain representations, underscoring the importance of input feature selection.
To improve latency and efficiency, alternative CNN architectures avoid spectrogram conversion. Spectrum Stitching [uvaydov2024stitching] uses a U-Net with a non-local block for direct I/Q processing, reducing computational overhead and speeding up inference. While this method eliminates feature extraction, it requires extensive training to generalize across signal variations, balancing accuracy and efficiency. In the context of O-RAN-based implementations, ChARM [9796985] provides a real-time, data-driven framework for dynamic spectrum sharing in O-RAN networks, enabling spectrum sensing, interference detection. However, its integration with O-RAN remains unclear, and its classifier handles a limited set of RF signals.
Our approach differs from previous works in multiple aspects. Unlike spectrogram-based classification ([zhang2021signal][amiciradiodl]) or hardware-efficient CNNs ([lees2019deep]), we process frequency-domain I/Q samples directly using a CNN trained on time-series data that performs multinomial classification, in contrast with the binary one of [zheng2020spectrum]. Before classification, we apply an Energy Peak Detector to isolate relevant signals and reduce dimensionality, optimizing both performance and efficiency. In terms of results, our method achieves 97.8% classification accuracy, significantly outperforming image-based classification ([amiciradiodl], 77.78%) and offering superior real-time adaptability compared to methods based on static spectrograms ([hauser2017signal], approx 90%). Furthermore, our spectrum classification model is the only one that has been tested within a real dApp framework, interacting directly with an operating gNB, unlike previous works that primarily rely on offline datasets or simulated environments or substantial modification of the O-RAN architecture [9796985].
@@ IV-A Dataset collection and preprocessing
We leverage and extend the dApp framework of [lacava2025dapps] to access and collect the I/Q samples in real-time. The dApp is designed with a sensing periodicity of approximately 8 ms, meaning it periodically provides data to LibIQ for labeling. Once collected, the I/Q samples are organized into complex input vectors, each containing real and imaginary components. Each vector consists of 1,536 I/Q samples, representing data collected from spectrum sensing. These vectors are stored in a binary file, with each file containing 100 input vectors.
We collected data from two different environments: one in an emulated environment using Colosseum [polese2024colosseum], and the other in a real environment using transmissions over the air. We use a 5G standalone (SA) deployment of one gNB with OpenAirInterface (OAI), operating at a center frequency of 3.6192 GHz with a 40 MHz bandwidth that onboard our dApp. As shown in Figure 1, for the OTA testbed we connect a Commercial Off-the-Shelf (COTS) One Plus Nord phone to the network to generate downlink traffic even though in this work we focus more on the external than the User Equipment (UE) performance. To generate external signals, we employed GNU Radio, the siggen utility of the uhd library for SDRs, and the SCOPE framework [bonati2021scope], each transmitting different signal types, as shown in Table I. The center frequency of the transmitted signals varied among 3.6042 GHz, 3.6142 GHz, 3.6242 GHz, and 3.6342 GHz, while the sampling rate remained fixed at 1 MHz throughout all transmissions.
@@
Sine $90$ $25$ 11 11 Radar Siggen
Triangular $100$ $25$ 11 11 Triangular GNU Radio
Square $100$ $25$ 11 11 Square GNU Radio
Uniform $80$ 2020 11 11 Jammer Siggen
LTE // 2020 // $0.9$ LTE SCOPE [bonati2021scope]
@@ V Results
Latency [ms] Time required for a single prediction
$2.23$
$\pm 0.48$
$1.13$
$\pm 0.04$
$0.92$
$\pm 0.11$
$0.90$
$\pm 0.05$